# Relativistic Roche problem for stars in precessing orbits around a spinning black hole

### v2

@@ III.5 Elliptic orbits
Figure 2: $\tilde{\Omega}^{2}$ as functions of $r_{0}/m$ for the first- and second-order resonances, along with the tidal disruption limit $\tilde{\Omega}^{2}_{\mathrm{crit}}$ as well as the expected values of $\tilde{\Omega}^{2}$ for typical values of BH-NS binaries and supermassive BH-white dwarf/ordinary star binaries. For both plots, we assumed $a/m=0.8$. The upper and lower panels show the results for $f=0$ and $f=1$.
@@ IV.1 Determination of resonances
However, during our research, we found that these matrices could have determinant 0 for particular equilibrium states. Furthermore, the terms in the right hand side were not in the image of the matrices. Thus the linear equations were not solvable for the particular cases in which the determinant vanishes. This would indicate that the ansatz used in these cases are inappropriate.
One has then a resonance. To solve the equations, a better ansatz would be: $\vec{\xi}=(\vec{\xi}_{0}+\vec{\xi}_{1}\tau)e^{i\omega_{\theta}\tau}$. In this case, the displacements would eventually not be infinitesimal. Thus our perturbative expansion breaks down at such points.
@@
In order to assess the relevancy of these resonances, we plot the corresponding values of $\tilde{\Omega}^{2}$ of these resonances, along with the values of $\tilde{\Omega}^{2}$ that are modeled by a neutron star, a white dwarf, and an ordinary star orbiting a variety of Kerr BHs. We estimate $\tilde{\Omega}^{2}$ using typical radius $R_{\mathrm{star}}$ and mass $M$ for these objects. We then have $\tilde{\Omega}^{2}$ as a function of $r_{0}$ as
@@
In Fig. 2, we plot the values of $\tilde{\Omega}^{2}$ for the resonances of $\omega=\omega_{\theta}$ and $2\omega_{\theta}$ as functions of $r_{0}/m$ together with the Roche $(f=0)$ and Roche-Riemann $(f=1)$ limits. We also plot the curves along hypothetical sequences of inspiraling BH-NS binaries in close orbits and white dwarf/ordinary stars orbiting supermassive BHs. We note that for the $\omega=0$ mode in the second-order displacement, the resonance appears near the Roche/Roche-Riemann limits. This corresponds to the onset of dynamical instability for the ellipsoid, similarly as in the classical Roche problem [2]. This figure shows that these resonances may destabilize the objects before reaching the Roche and Roche-Riemann limits. While we show figures only for one value of $a/m(=0.8)$, we observe the same behavior for any non-zero value of $a/m$.
@@
In our present context, the displacement associated with an oscillatory mode gets excited by the variation of the $\theta$-coordinate of the orbit. It will then grow linearly and eventually, it will not be infinitesimal anymore. Figures 3 and 4 show the displacement vectors associated with the excited modes, assuming $a/m=0.8$. Again the spin parameter does not have an important impact on the qualitative aspect of the resonant displacement.
From Fig. 3, we can see that the resonance in the first-order displacement will mostly imply a precession of the 3-axis. As such, this mode is quite reminiscent of the transverse-shear mode of the MacLaurin spheroids discussed in [2]. By the non-linear amplification of this mode, matter may be ejected from the surface to the outside of the star, resulting possibly in the modification of the precessing orbit.
@@ IV.2 Analytical resolution of the first-order resonance
$\begin{split}\det(M_{ij}^{(1)})=&-16\pi\rho B_{23}(1+\lambda)\Omega_{0}^{4}\\ &\times\Big{[}-4\pi\rho B_{13}\lambda+(3+\lambda)(1+\lambda)\Omega_{0}^{2}\Big{]},\end{split}$ (101)
@@ V Discussion
To our knowledge, these results have never been reported before. A major counterpoint to our analysis is its apparent reliance on some questionable assumptions, like the incompressibility hypothesis. However, as the general mechanism for the resonance holds true, we expect resonances even with more realistic equations of state. Indeed, whatever the equation of state is, we have characteristic modes for our star. We may then observe a resonance between these modes and the oscillation present in the tidal tensor. Thus, while the equations can change, the resulting phenomenon should not qualitatively. Of course, this explanation is just tentative and should be confirmed by a more complete analysis of the question.
It is possible that this resonance might provide an explanation as to why certain objects end up in a circular equatorial orbit around a Kerr BH. To justify this, one would have to inquire about the ultimate fate of the star. As our equations are not valid near the resonance, one would have to change our method. One could use different analytical tools or hydrodynamics simulations to investigate this issue. We leave this open to future work.
@@ Appendix A Formula for second-order equations
$\begin{split}\Omega^{(2)}&=a^{2}\cos^{2}\theta\frac{\sqrt{m}(r_{0}^{2}-3mr_{0}+2a\sqrt{mr_{0}})}{r_{0}^{7/2}(\sqrt{mr_{0}}-a)^{2}}\\ &+\mathcal{C}\left[2K_{1}\Omega_{0}+\frac{\sqrt{K_{0}}}{r_{0}^{2}}\left(\frac{E_{1}(r_{0}^{2}+a^{2})-aL_{1}}{r_{0}^{2}+K_{0}}-\frac{K_{1}(E_{0}(r_{0}^{2}+a^{2})-aL_{0})}{(r_{0}^{2}+K_{0})^{2}}+\frac{a(L_{1}-aE_{1})}{K_{0}}-\frac{aK_{1}(L_{0}-aE_{0})}{K_{0}^{2}}\right)\right]\\ &=a^{2}\cos^{2}\theta\Omega_{a}+\mathcal{C}\Omega_{\mathcal{C}}.\end{split}$ (109)

### v3

@@ III.5 Elliptic orbits
Figure 2: $\tilde{\Omega}^{2}$ as functions of $r_{0}/m$ for the first- and second-order resonances, along with the tidal disruption limit $\tilde{\Omega}^{2}_{\mathrm{crit}}$ as well as the expected values of $\tilde{\Omega}^{2}$ for typical values of BH-NS binaries and supermassive BH-white dwarf/ordinary star binaries. For both plots, we assumed $a/m=0.8$ and prograde orbits. The upper and lower panels show the results for $f=0$ and $f=1$.
Figure 3: The same as Fig. 2 but for $a/m=0.5$ (upper and middle panels) and for $a/m=0.9$ (lower panel). The upper panel shows the case of retrograde orbits, while the middle and lower panels prograde orbits.
@@ IV.1 Determination of resonances
However, during our research, we found that these matrices could have determinant zero for particular equilibrium states. Furthermore, the terms in the right hand side were not in the image of the matrices. Thus the linear equations were not solvable for the particular cases in which the determinant vanishes, implyimg the presence of a resonance. This indicate that the ansatz used in these cases are inappropriate. To solve the equations at the resonance point, a better ansatz would be: $\vec{\xi}=(\vec{\xi}_{0}+\vec{\xi}_{1}\tau)e^{i\omega_{\theta}\tau}$. In this case, the displacements would eventually not be infinitesimal. Thus our perturbative expansion breaks down at such points.
@@
In order to assess the relevancy of these resonances, we plot the corresponding values of $\tilde{\Omega}^{2}$ of these resonances, along with the values of $\tilde{\Omega}^{2}$ that are modeled by an NS, a white dwarf, and an ordinary star orbiting a variety of Kerr BHs. We estimate $\tilde{\Omega}^{2}$ using typical radius $R_{\mathrm{star}}$ and mass $M$ for these objects. We then have $\tilde{\Omega}^{2}$ as a function of $r_{0}$ as
@@
In Fig. 2, we plot the values of $\tilde{\Omega}^{2}$ for the resonances of $\omega=\omega_{\theta}$ and $2\omega_{\theta}$ as functions of $r_{0}/m$ together with the Roche $(f=0)$ and Roche-Riemann $(f=1)$ limits, for $a/m=0.8$. We also plot the curves along hypothetical sequences of inspiraling BH-NS binaries in close orbits and white dwarf/ordinary stars orbiting supermassive BHs. We note that for the $\omega=0$ mode in the second-order displacement, the resonance appears near the Roche/Roche-Riemann limits. This corresponds to the onset of dynamical instability for the ellipsoid, similarly as in the classical Roche problem [2]. This figure shows that these resonances may destabilize the objects before reaching the Roche and Roche-Riemann limits.
In Fig. 3, we reproduce the same figure for $a/m=0.5$ and $a/m=0.9$ with $f=0$. We find that the spin parameter $a/m$ does not have much of an impact on the relative position of the resonances to the Roche limit. As in [8], we see that the Roche limit gets further away and with it the resonances, as $a/m$ decreases.
In the upper panel of Fig. 3, we produced the curves in the case of retrograde orbits. While until now we always presented the results for prograde orbits, they are easily extended to retrograde orbits. One only has to take $a/m<0$. The observation made before can then also be extended. Compared to positive values of $a/m$, we observe that the resonances happen for larger values of $r_{0}/m$. However their relative positions to the Roche limit are the same.
@@
In our present context, the displacement associated with an oscillatory mode gets excited by the variation of the $\theta$-coordinate of the orbit. It will then grow linearly and eventually, it will not be infinitesimal anymore. Figures 4 and 5 show the displacement vectors associated with the excited modes, assuming $a/m=0.8$. Again the spin parameter does not have an important impact on the qualitative aspect of the resonant displacement.
Fig. 4 shows that the resonance in the first-order displacement will mostly imply a precession of the 3-axis. As such, this mode is quite reminiscent of the transverse-shear mode of the MacLaurin spheroids discussed in [2]. By the non-linear amplification of this mode, matter may be ejected from the surface to the outside of the star, resulting possibly in the modification of the precessing orbit as well as the stellar profile. For white dwarfs, the stellar radius increases with the mass ejection. Thus, this resonance could trigger the tidal disruption.
@@
We note that in our analysis we neglected several effects, which can modify our results quantitatively. Since we neglected the general relativistic effect for the self-gravity of the NS, an error of order $Gm_{\mathrm{NS}}/c^{2}R_{\mathrm{star}}$ may be expected on the condition of the tidal disruption. Also we neglected the NS’s gravity on the orbital motion, so that an error of order $M/m$ may be also expected. However, these dimensionless parameters are supposed to be of order 0.1 in this paper. Hence, the results obtained in this paper is unlikely to be significantly changed even if we take into account these effects.
@@ IV.2 Analytical resolution of the first-order resonance
$\det(M_{ij}^{(1)})$ $=$ $-16\pi\rho B_{23}(1+\lambda)\Omega_{0}^{4}$ (102)
×$\times\Big{[}-4\pi\rho B_{13}\lambda+(3+\lambda)(1+\lambda)\Omega_{0}^{2}\Big{]},~{}~{}~{}$
@@ V Discussion
As such, the instability found is unlike that found at the Roche limit. Indeed, we do not provide a correction to this limit as in [9]. Instead, we provide a resonance which arises due to the dynamical nature of our analysis. Thus our results and their interpretation differ from [9].
@@
The second-order resonance can be relevant for BH-NS binaries just prior to the significant tidal deformation of the NS. In the presence of a precession, this resonance is excited during the late inspiral phase, and thus, the orbital energy is partly transported to the oscillation energy of the NS. This can accelerate the orbital evolution of the system, leading to a gravitational-wave phase shift, in addition to the phase shift related to the usual tidal deformation [20, 21]. Since the second-order resonance can be induced only at an orbit close to the tidal disruption limit, it may not be easy to distinguish the resonant effect from the usual tidal deformation effect, and thus, we may overestimate the tidal deformation effect if we do not properly take into account the resonant effect for precessing binaries.
To our knowledge, these results have never been reported before. A major counterpoint to our analysis is its apparent reliance on some questionable assumptions, like the incompressibility hypothesis. However, as the general mechanism for the resonance holds true, we expect resonances even with more realistic equations of state. Indeed, whatever the equation of state is, we have characteristic modes for our star. We may then observe a resonance between these modes and the oscillation present in the tidal tensor. Thus, while the equations can change, the resulting phenomenon should not qualitatively. Of course, this explanation is just tentative and should be confirmed by a more complete analysis of the question. We plan to show the results of our approximate analysis for the compressible equations of state in a follow-up paper (Stockinger and Shibata, in preparation).
As our equations are not valid near the resonance, one would have to change our method to fully understand the effects on the orbital change by the resonance. For this purpose, performing hydrodynamics simulations is a robust approach. We are currently working on simulations of these types of systems (Lam et al., in preparation).
@@ Appendix A Formula for second-order equations
$\Omega^{(2)}$ $=$ $a^{2}\cos^{2}\theta\frac{\sqrt{m}(r_{0}^{2}-3mr_{0}+2a\sqrt{mr_{0}})}{r_{0}^{7/2}(\sqrt{mr_{0}}-a)^{2}}$ (110)
$+\mathcal{C}\left[2K_{1}\Omega_{0}+\frac{\sqrt{K_{0}}}{r_{0}^{2}}\left(\frac{E_{1}(r_{0}^{2}+a^{2})-aL_{1}}{r_{0}^{2}+K_{0}}-\frac{K_{1}(E_{0}(r_{0}^{2}+a^{2})-aL_{0})}{(r_{0}^{2}+K_{0})^{2}}+\frac{a(L_{1}-aE_{1})}{K_{0}}-\frac{aK_{1}(L_{0}-aE_{0})}{K_{0}^{2}}\right)\right]$
$=$ $a^{2}\cos^{2}\theta\Omega_{a}+\mathcal{C}\Omega_{\mathcal{C}}.$
@@
Appendix B Quadratic contributions to the second order
The quadratic contributions to the second order can be divided in three parts
$Q_{ij}(V_{k;l})=\frac{5}{M}(Q_{ij}^{\text{grav}}(V_{k;l})+Q_{ij}^{\text{stat}}(V_{k;l})+Q_{ij}^{\text{dyn}}(V_{k;l})),$
where
$Q_{ij}^{\text{stat}}(V_{k;l})=\sum_{k=1}^{3}\sum_{l=1}^{3}(\Omega_{0}^{2}(\delta_{ik}-\delta_{3k})-\tilde{C}^{(0)}_{ik})\frac{V_{k;l}V_{j;l}}{a_{l}^{2}},$
$Q_{ij}^{\text{dyn}}(V_{k;l})$ $=$ $\sum_{k=1}^{3}\bigg{(}-\frac{\ddot{V}_{i;k}V_{j;k}}{a_{k}^{2}}+2\sum_{l=1}^{3}\big{(}Q_{jl}\dot{V}_{i;k}$
$-(Q^{2})_{jl}V_{i;k}+(Q^{2})_{il}V_{j;k}\big{)}\frac{V_{l;k}}{a_{k}^{2}}$
$-2\sum_{l=1}^{3}\epsilon_{il3}\Omega_{0}\big{(}\dot{V}_{l;k}\frac{V_{j;k}}{a_{k}^{2}}$
$+\sum_{m=1}^{3}(Q_{lm}V_{j;k}-Q_{jm}V_{l;k})\frac{V_{m;k}}{a_{k}^{2}}\big{)}\bigg{)}$
and finally
$Q_{ij}^{\text{grav}}(V_{k;l})$ $=$ $\sum_{l=1}^{3}\big{(}\frac{\mathcal{A}_{i}V_{i;l}V_{j;l}}{a_{l}^{2}}+\frac{\mathcal{A}_{ij}V_{i;l}V_{j;l}a_{j}^{2}}{a_{l}^{2}}+\mathcal{A}_{ilj}V_{il}V_{lj}a_{j}^{2}$ (121)
$-V_{il}\mathcal{A}_{il}V_{lj}+V_{ll}\mathcal{A}_{il}V_{ij}+\frac{V_{ll}V_{ij}\mathcal{A}_{ijl}a_{j}^{2}}{2}$
$+\delta_{ij}\sum_{k=1}^{3}(-\frac{\mathcal{A}_{i}V_{l;k}^{2}a_{j}^{2}}{a_{l}^{2}a_{k}^{2}}-\frac{V_{lk}^{2}\mathcal{A}_{lki}a_{j}^{2}}{4}$
$-\frac{\mathcal{A}_{li}V_{l;k}^{2}a_{j}^{2}}{2a_{k}^{2}}+\frac{\mathcal{A}_{lki}V_{ll}V_{kk}a_{j}^{2}}{8})\big{)}.$
When we consider a displacement $V_{i;j}$ whose non-zero terms have indices $(i,j)=(1,3),(3,1),(2,3),(3,2)$, the quadratic contributions are non-zero only for the indices even on the component 3.